# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

### v2

@@ Results.
AWQ. A major strength of our algorithm is its versatility. It seamlessly replaces GPTQ in any quantization technique that relies on it. To illustrate this, we focus on the AWQ algorithm. We demonstrate that our algorithm, when layered on top of AWQ, surpasses the performance of GPTQ layered on top of AWQ. Table 3 presents the result from this experiment. It can be seen that BCD always outperforms GPTQ, for both per-channel and group quantization  [^2] The performance of AWQ on PaLM2-Bison is much worse than not performing AWQ. Hence, in Table 3, we only present results for PaLM2-Otter..
Runtime.
The runtimes of various techniques for quantizing PaLM2-Otter and PaLM2-Bison are presented in Table 6. To be precise, the table presents time taken to quantize all the FFN, attention layers in the model, for per-channel quantization. For group quantization runtimes, please refer to Appendix E. It can be seen that CD and GPTQ have comparable runtimes. While BCD takes longer than CD, the majority of its time is spent quantizing the $2^{nd}$ FFN layer, where the quantization dimension is equivalent to the model’s hidden dimension. So, in practice, to speed up the entire process, one could rely on CD for quantizing FFN2 and use BCD for quantizing the rest of the weight matrices.
Table 4: Table presents pretraining perplexity numbers for GPTQ, CD, BCD for INT3, INT4 quantization of both FFN and attention weights.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w16a16 - 7.948 5.984 5.298
w3a16 OWC $41.751$ $54.499$ $6.446$
GPTQ $24.054$ $\mathbf{11.384}$ $5.801$
CD $19.692$ $11.542$ $5.770$
BCD(k=2) $\mathbf{19.312}$ $11.392$ $\mathbf{5.762}$
w3a16g128 OWC $20.301$ $9.785$ $5.973$
GPTQ $15.042$ $7.501$ $5.698$
CD $14.757$ $7.454$ $5.679$
BCD(k=2) $14.691$ $\mathbf{7.438}$ $5.673$
OWC-CD $16.268$ $7.617$ $5.746$
OWC-CD + CD $14.285$ $7.466$ $5.675$
OWC-CD + BCD(k=2) $\mathbf{14.242}$ $7.458$ $\mathbf{5.670}$
w4a16 OWC $10.683$ $7.194$ $5.508$
GPTQ $10.005$ $6.645$ $5.425$
CD $9.991$ $6.627$ $5.415$
BCD(k=2) $\mathbf{9.965}$ $\mathbf{6.621}$ $\mathbf{5.413}$
w4a16g128 OWC $9.402$ $6.415$ $5.420$
GPTQ $9.160$ $6.241$ $5.381$
CD $9.124$ $6.226$ $5.377$
BCD(k=2) $9.120$ $\mathbf{6.225}$ $5.376$
OWC-CD $9.241$ $6.261$ $5.387$
OWC-CD + CD $9.094$ $6.244$ $5.375$
OWC-CD + BCD(k=2) $\mathbf{9.093}$ $6.242$ $\mathbf{5.374}$
Table 5: Table presents the perplexity, and downstream evaluations (1-shot) for GPTQ, CD, BCD for INT2 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Method Epochs PaLM2-Otter PaLM2-Bison
w2a16g128 Perplexity Generation Rank Avg. Perplexity Generation Rank Avg.
GPTQ - $10.816$ $26.98$ $67.14$ $51.07$ $7.230$ $39.32$ $74.19$ $60.24$
CD $1$ $9.917$ $29.87$ $66.76$ $52.00$ $7.123$ $38.73$ $73.56$ $59.63$
BCD(k=2) $1$ $9.822$ $30.03$ $66.82$ $52.11$ $7.094$ $38.85$ $73.99$ $59.93$
$2$ $9.760$ $30.09$ $\mathbf{67.45}$ $\mathbf{52.51}$ $\mathbf{7.088}$ $\mathbf{39.35}$ $74.10$ $60.20$
$3$ $9.719$ $\mathbf{30.19}$ $67.09$ $52.33$ $7.094$ $39.30$ $\mathbf{74.24}$ $\mathbf{60.26}$
BCD(k=4) $1$ $\mathbf{9.708}$ $30.11$ $67.09$ $52.30$ $7.099$ $38.98$ $74.20$ $60.11$
$2$ $9.752$ $30.00$ $66.70$ $52.02$ $7.099$ $39.09$ $74.08$ $60.08$
$3$ $9.749$ $30.07$ $66.93$ $52.19$ $7.099$ $39.05$ $74.10$ $60.08$
Table 6: Table presents the per-channel quantization runtime (in minutes) of CD, BCD, GPTQ (each run for $d_{\text{in}}$ iterations) on PaLM2-Otter and PaLM2-Bison.
Config Method PaLM2-Otter PaLM2-Bison
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
FFN runtime
(in minutes)
Attn. runtime
(in minutes)
w3a16 GPTQ $0.90$m $0.72$m $6.42$m $0.63$m
CD $2.94$m $0.49$m $16.13$m $0.43$m
BCD(k=2) $14.03$m $2.08$m $82.13$m $1.81$m
w4a16 GPTQ $0.90$m $0.64$m $6.60$m $0.56$m
CD $3.59$m $0.59$m $21.19$m $0.52$m
BCD(k=2) $34.85$m $4.72$m $235.14$m $4.13$m
@@ 5 Conclusion and Future Work
In this work, we developed a coordinate descent framework (CDQuant) for quantization of LLMs. CDQuant is a simple and effective alternative to GPTQ, that consistently outperformed it on PaLM2 models. The simplicity of our algorithm makes it a seamless substitute for GPTQ in various algorithmic contexts where GPTQ currently functions as a sub-routine. Our future work aims to further improve the performance of CDQuant. In particular, we aim to speed up our BCD algorithm, and make it as fast as CD. Furthermore, we will focus on developing layer-wise loss functions that are more closely aligned with end-to-end loss, thereby reducing the performance gap between full-precision and quantized models. Finally, we will also explore integrating CDQuant with QuIP and FrameQuant to evaluate potential performance gains for extreme low-bit quantization.
@@ Appendix B Limitations
  • Due to computational reasons, we couldn’t run experiments where we replace GPTQ component with CDQuant in algorithms such as QuIP, FrameQuant.
Appendix C CDQuant for sub-channel quantization
@@
Appendix D Detailed Downstream Evaluation Results
Table 7: Table presents downstream evaluation (one-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Bison
@@ Appendix D Detailed Downstream Evaluation Results
w16a16 - $27.84$ $75.19$ $77.42$ $23.87$ $59.04$ $84.64$ $88.87$ $82.50$ $82.43$ $78.22$
w3a16 OWC $22.16$ $75.84$ $70.31$ $21.75$ $57.34$ $83.50$ $86.94$ $79.60$ $81.12$ $78.06$
GPTQ $24.29$ $74.65$ $72.29$ $22.98$ $58.70$ $85.02$ $87.58$ $80.81$ $81.72$ $77.66$
CD $23.60$ $74.92$ $72.25$ $23.13$ $58.53$ $84.68$ $88.20$ $80.60$ $81.61$ $77.74$
BCD(k=2) $23.24$ $74.78$ $72.06$ $22.24$ $58.62$ $84.55$ $88.04$ $80.69$ $81.83$ $78.22$
w3a16g128 OWC $23.88$ $73.59$ $73.31$ $21.56$ $57.85$ $83.84$ $88.23$ $80.67$ $80.79$ $78.69$
GPTQ $25.32$ $73.96$ $74.10$ $23.23$ $58.02$ $84.39$ $88.50$ $80.67$ $81.83$ $78.06$
CD $25.24$ $74.16$ $74.11$ $23.18$ $57.51$ $83.75$ $88.38$ $81.15$ $82.15$ $77.98$
BCD(k=2) $24.93$ $74.05$ $73.90$ $23.18$ $56.74$ $84.05$ $88.26$ $81.06$ $82.43$ $78.53$
OWC-CD $24.60$ $74.03$ $73.80$ $21.75$ $57.17$ $84.18$ $88.2$ $80.97$ $81.94$ $77.27$
OWC-CD + CD $25.12$ $74.37$ $73.99$ $23.38$ $58.36$ $85.44$ $88.65$ $80.93$ $81.77$ $77.98$
OWC-CD + BCD(k=2) $25.32$ $73.54$ $74.30$ $23.18$ $57.34$ $83.96$ $88.59$ $81.01$ $81.39$ $77.51$
w4a16 OWC $26.57$ $74.66$ $75.84$ $23.13$ $58.28$ $83.42$ $88.75$ $81.39$ $81.66$ $78.85$
GPTQ $27.06$ $74.87$ $76.45$ $22.74$ $58.87$ $84.60$ $88.53$ $81.72$ $82.48$ $78.69$
CD $26.79$ $74.40$ $76.35$ $22.83$ $59.22$ $84.68$ $88.99$ $81.60$ $82.10$ $78.69$
BCD(k=2) $27.06$ $74.54$ $76.28$ $23.13$ $59.81$ $84.76$ $88.56$ $81.70$ $82.26$ $77.74$
w4a16g128 OWC $27.04$ $75.37$ $76.42$ $23.13$ $59.73$ $84.97$ $88.87$ $81.94$ $82.32$ $78.45$
GPTQ $27.12$ $75.63$ $76.42$ $23.38$ $58.45$ $84.76$ $89.05$ $82.01$ $82.15$ $78.85$
CD $27.31$ $75.52$ $76.45$ $23.62$ $59.04$ $85.02$ $88.50$ $82.04$ $82.54$ $78.69$
BCD(k=2) $27.37$ $75.44$ $76.92$ $23.62$ $58.87$ $85.35$ $88.72$ $81.98$ $82.97$ $79.48$
OWC-CD $27.31$ $75.70$ $76.64$ $22.93$ $58.36$ $84.64$ $88.56$ $81.98$ $82.59$ $78.14$
OWC-CD + CD $27.01$ $75.52$ $76.69$ $23.28$ $58.79$ $85.10$ $88.56$ $81.98$ $82.15$ $78.69$
OWC-CD + BCD(k=2) $26.98$ $75.42$ $76.84$ $23.33$ $58.79$ $85.35$ $88.47$ $81.93$ $82.32$ $78.14$
Table 8: Table presents downstream evaluation (one-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.

### v3

@@ 4.2 Versatility of CDQuant
A major strength of our algorithm is its versatility. It seamlessly replaces GPTQ in any quantization technique that relies on it. To illustrate this, we focus on the AWQ, QuIP and FrameQuant (the latter two are state-of-the-art PTQ techniques). We demonstrate that our algorithm, when layered on top of the aforementioned algorithms, surpasses the performance of GPTQ layered on top of them. Table 3 presents the results for QuIP, FrameQuant, and Table 4 presents the results for AWQ. It can be seen that both CD and BCD provide $\sim 5\%$ boost in performance for QuIP and FrameQuant. For AWQ, we observe that CD, BCD always outperform GPTQ, for both per-channel and sub-channel quantization (the performance of AWQ on PaLM2-Bison is much worse than not performing AWQ. Hence, in Table 4, we only present results for PaLM2-Otter from the PaLM2 family of models).
Table 7: Runtime (in minutes using 8 H100s) comparison of CDQuant, GPTQ for Gemma-2 27B. Note that FFN1 and FFN1-Gate have the same runtimes, hence we report only FFN1’s runtime.
Config Method
FFN1
FFN2
Attn.
w3a16 GPTQ $0.3$ $7.37$ $1.18$
CD $1.23$ $14.1$ $0.51$
BCD(k=2) $6.03$ $54.94$ $2.11$
w4a16 GPTQ $0.3$ $7.35$ $1.18$
CD $1.65$ $17.87$ $0.64$
BCD(k=2) $14.54$ $133.34$ $4.64$
Table 8: Table presents the perplexity evaluations for INT3 quantization with BCD wherein the FFN2 may be quantized either with CD or BCD, while rest of the FFN parameters are quantized with BCD.
Config Method Gemma-1 7B Gemma-2 9B
w16a16 $10.384$ $10.683$
w3a16 CD, FFN2 $19.709$ $11.288$
BCD(k=2), FFN2 $18.552$ $11.284$
w3a16g128 CD, FFN2 $13.512$ $11.179$
BCD(k=2), FFN2 $13.501$ $11.180$
5 Runtime
Table 8 shows the runtime of GPTQ, CD and BCD for quantization of FFN1, FFN2, and attention weights of Gemma-2 27B model using 8 H100s (see Appendix F.4 for 1 GPU runtime). For attention weight quantization, CD is comparable to GPTQ in speed, while BCD is 2×2$\times$2 × slower. For FFN1 (FFN2) quantization, CD is 5×5$\times$5 × (2×2$\times$2 ×) slower than GPTQ, whereas BCD is an order of magnitude slower than GPTQ. It can be seen that FFN2 quantization is the most time-consuming step (because the quantization axis is of size $d_{\text{hidden}}$ which is much larger than the quantization axis of other layers). Consequently, CD (BCD) is twice (order of magnitude) as slow as GPTQ for whole-model quantization. We now provide two simple strategies to speed up both CD and BCD.
Speeding up BCD. To speed up BCD, we quantize FFN2 with CD, and rest of the layers with BCD (where the quantization axis is of size $d_{\text{model}}$). Results for this setting can be found in Table 8. It can be seen that quantizing FFN2 with CD leads to negligible to no drop in performance especially for the larger, and better Gemma-2 9B. With this setup, BCD’s runtime is reduced significantly and is comparable to that of CD.
Speeding up CD. We find that CD converges to an optimal solution in much fewer iterations than $d_{\text{in}}$ iterations used in our experiments. Based on this finding, we run CD for lesser number of iterations. As shown in Table 22 in Appendix, even with a drastically reduced iteration count (e.g., $d_{\text{in}}/8$), CD maintains near-optimal performance with negligible quality loss, surpassing the performance of GPTQ. This reduction in iterations allows CD to achieve a runtime faster than/comparable to GPTQ.
@@ 6 Conclusion and Future Work
In this work, we developed a coordinate descent framework (CDQuant) for quantization of LLMs. CDQuant is a simple and effective alternative to GPTQ, that consistently outperformed it on PaLM2 models. The simplicity of our algorithm makes it a seamless substitute for GPTQ in various algorithmic contexts where GPTQ currently functions as a sub-routine. Our future work aims to further improve the performance of CDQuant. In particular, we would like to use ideas from  CDQuant to improve Quantized Aware Training (QAT). Furthermore, we will focus on developing layer-wise loss functions that are more closely aligned with end-to-end loss, thereby reducing the performance gap between full-precision and quantized models.
@@ Appendix B Limitations
Appendix C Related Work
In this section, we provide a more thorough comparison of our technique with OBQ and GPTQ.
Detailed comparison with OBQ.
At a high level, both OBQ, and our technique use greedy strategies to quantize the weights. However, OBQ is extremely slow and doesn’t even scale to models with a few billion parameters. The primary reason for this unreasonable slowness is because of a more complicated step involved in OBQ which requires updating all the unquantized coordinates in each step (please refer to equation 2 in Frantar & Alistarh (2022)). Implementing this step in GPUs is extremely slow because of the sheer number of gathers/scatters that need to be performed at each step. In contrast, CDQuant doesn’t have this step. At each step of coordinate descent, we only update the single coordinate that leads to the biggest drop in performance. As a result, CDQuant is orders of magnitude faster than OBQ. To quantify this last statement, note that GPTQ paper only ran OBQ for models with a few hundred millions of parameters and showed that OBQ is almost 120x slower than GPTQ (note that this difference will only grow with larger models because OBQ will require even more gathers/scatters). In fact, in the OBQ paper (Frantar & Alistarh, 2022), in Table 6, OBQ takes 65 minutes to quantize a ResNet50 model with 25M parameters. That being said, scaling it to models with billions of parameters seems extremely difficult.
Detailed comparison with GPTQ.
GPTQ, as discussed in Section 2, is a heuristic designed to accelerate OBQ by replacing its greedy strategy with a single, predetermined cycle through all coordinates. While this approach improves speed, it often sacrifices accuracy. Although multiple cycles could potentially mitigate this performance drop, our experiments revealed no significant improvement in quality. In contrast, we propose a simple yet effective greedy coordinate descent strategy that achieves superior accuracy while maintaining the computational efficiency of GPTQ.
Appendix D Comparison to QuantEase
In this section, we compare our per-channel quantization results with those of QuantEase, a parallel study to ours. It’s important to note that QuantEase does not have a publicly available implementation. So, we implemented the cyclic coordinate descent strategy used by QuantEase, and used the same initialization and regularization strength as our algorithms (although the QuantEase paper doesn’t provide these details). We then ran QuantEase for the recommended number of iterations specified in the paper ($20$ epochs or $T=20d_{\text{in}}$ iterations). The findings are presented in Table 9. On PaLM2-Gecko, PaLM2-Otter, and PaLM2-Bison, both these approaches have similar performance. These results collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ, especially for low bit quantization.
Table 9: Comparison of QuantEase with CD and BCD for FFN quantization.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
w3a16 GPTQ $11.347$ $7.176$ $5.774$
QuantEase (epochs=20) $10.731$ $6.996$ $5.741$
CD $10.920$ $7.002$ $5.739$
BCD(k=2) $10.898$ $6.979$ $5.733$
w4a16 GPTQ $8.764$ $6.249$ $5.417$
QuantEase (epochs=20) $8.670$ $6.197$ $5.408$
CD $8.694$ $6.195$ $5.407$
BCD(k=2) $8.691$ $6.192$ $5.405$
Appendix E CDQuant for sub-channel quantization
@@
Appendix F Additional Experimental Results
This section presents additional experimental results:
  1. 1.
    In Appendix F.1, we present the effects of a larger block size and multiple epochs on BCD.
  2. 2.
    In Appendix F.2, we present a detailed breakdown of downstream evaluation results on various datasets.
  3. 3.
    In Appendix F.3, we present results for the setting where we quantize both attention and FFN layers.
  4. 4.
    In Appendix F.4, we present runtime numbers for CD and BCD algorithms
F.1 Additional ablations for BCD
We investigate the effects of a large block size and multiple epochs on BCD, where each epoch corresponds to $d_{\text{in}}$ iterations. Results for this experiment are presented in Table 10. For BCD with $k=2$, additional epochs enhance performance. However, for BCD with $k=4$, multiple epochs appear to lead to overfitting, indicating a need for stronger regularization.
Table 10: Table presents the perplexity evaluations BCD for INT2 quantization of FFN weights when run with a larger block size and multiple epochs.
Method Epochs Gemma-1 7B Gemma-2 9B Gemma-2 27B PaLM2-Otter PaLM2-Bison
w16a16 $10.384$ $10.683$ $8.682$ $5.984$ $5.298$
w2a16g128
GPTQ - $375.153$ $13.785$ $12.181$ $10.816$ $7.230$
CD $1$ $75.55$ $13.709$ $11.966$ $9.917$ $7.123$
BCD(k=2) $1$ $\mathbf{68.732}$ $\mathbf{13.662}$ $\mathbf{11.873}$ $\mathbf{9.822}$ $\mathbf{7.094}$
$2$ $67.766$ $13.665$ $11.888$ $9.760$ $\mathbf{7.088}$
$3$ $72.330$ $13.674$ $11.878$ $9.719$ $7.094$
BCD(k=4) $1$ $68.785$ $\mathbf{13.645}$ $\mathbf{11.857}$ $\mathbf{9.708}$ $7.099$
$2$ $67.390$ $13.640$ $11.867$ $9.752$ $7.099$
$3$ $67.590$ $13.664$ $11.869$ $9.749$ $7.099$
F.2 Detailed Downstream Evaluation Results
Config Method PaLM2-Gecko
@@ F.2 Detailed Downstream Evaluation Results
w16a16 - $4.68$ $43.17$ $27.46$ $5.07$ $35.24$ $65.66$ $60.24$ $60.69$ $74.7$ $61.48$
w3a16 OWC $2.55$ $47.57$ $9.81$ $1.87$ $29.69$ $56.73$ $64.8$ $53.28$ $69.59$ $60.3$
GPTQ $3.35$ $44.13$ $3.97$ $0.89$ $30.2$ $62.12$ $63.79$ $56.33$ $71.22$ $57.85$
CD $3.55$ $45.71$ $11.32$ $2.71$ $30.63$ $60.56$ $66.45$ $57.2$ $71.87$ $62.43$
BCD(k=2) $3.63$ $47.04$ $11.22$ $2.76$ $30.29$ $60.52$ $65.32$ $57.33$ $71.82$ $61.8$
w3a16g128 OWC $3.68$ $55.38$ $12.5$ $2.61$ $31.66$ $60.14$ $66.12$ $53.19$ $71.71$ $58.88$
GPTQ $4.02$ $50.04$ $14.64$ $3.15$ $32.42$ $61.91$ $66.64$ $56.05$ $73.01$ $59.75$
CD $3.38$ $52.01$ $9.62$ $2.21$ $31.23$ $63.13$ $65.9$ $55.98$ $72.47$ $59.91$
BCD(k=2) $3.1$ $50.95$ $9.36$ $1.92$ $31.14$ $62.25$ $65.41$ $56.16$ $72.91$ $58.96$
OWC-CD $3.6$ $51.69$ $9.33$ $2.31$ $31.83$ $61.74$ $65.02$ $56.23$ $72.2$ $59.43$
OWC-CD + CD $3.82$ $50.39$ $9.5$ $1.77$ $32.51$ $62.25$ $66.33$ $56.91$ $71.27$ $60.06$
w4a16 OWC $4.02$ $46.96$ $10.63$ $2.66$ $33.19$ $61.99$ $66.7$ $59.26$ $73.67$ $61.72$
GPTQ $4.46$ $48.69$ $18.84$ $3.1$ $34.47$ $62.96$ $64.77$ $60.09$ $73.61$ $61.09$
CD $4.13$ $48.88$ $16.64$ $3.15$ $33.53$ $63.89$ $66.79$ $60.31$ $73.78$ $61.4$
BCD(k=2) $4.24$ $48.31$ $15.73$ $3.3$ $33.28$ $63.34$ $66.39$ $60.26$ $73.56$ $61.01$
w4a16g128 OWC $4.99$ $46.1$ $20.62$ $4.28$ $34.73$ $64.48$ $67.83$ $59.91$ $74.32$ $61.25$
GPTQ $5.18$ $49.79$ $18.17$ $3.15$ $35.15$ $64.6$ $66.24$ $60.21$ $74.32$ $61.88$
CD $5.73$ $46.96$ $24.17$ $4.53$ $36.18$ $66.04$ $67.8$ $60.7$ $73.67$ $61.25$
BCD(k=2) $5.79$ $47.28$ $24.11$ $4.58$ $36.01$ $65.7$ $67.06$ $60.73$ $73.61$ $61.01$
OWC-CD $5.54$ $47.24$ $21.71$ $4.72$ $34.3$ $64.73$ $68.01$ $60.72$ $73.56$ $61.01$
OWC-CD + CD $5.62$ $47.6$ $24$ $4.68$ $35.49$ $66.16$ $67.06$ $60.78$ $74.1$ $61.48$